OpenStudy (waheguru):

3a-2(5a+1)+11=4(5-3a)-1

OpenStudy (waheguru):

\[3a-2(5a+1)+11=4(5-3a)-1 \]

OpenStudy (anonymous):

Start by distribution\[3a-2(5a)+2(1)+11=4(5)-4(3a)-1\]

OpenStudy (anonymous):

the third term was supposed to be negative

OpenStudy (waheguru):

I have a question about one step \[3a-10-2+11=20-12a-1\] \[3a-1=19-12a \] If I bring the -12 over will it be positive or nrgative

OpenStudy (waheguru):

For the second setp

OpenStudy (waheguru):

3a−1=19−12a it should be 3a+12a-1=19 right?

OpenStudy (anonymous):

You left out the "a" in 10a in your post up there^

OpenStudy (waheguru):

Sry, but my question is will the 12 be postive or negative if i bring it over Do you guys understant my quesiton >

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

It will be positive. Technically what you are doing is adding +12 to both sides. This causes the -12 on the right side to disappear, and +12 to appear on the left

OpenStudy (waheguru):

Oops, the problem was the fact I for to add the a in 10 a Yea, I was right thanks

OpenStudy (anonymous):

It will be positive..

OpenStudy (waheguru):

Thanks Guys

hero (hero):

\(3a-2(5a+1)+11=4(5-3a)-1 \\ 3a - 10a - 2 + 11= 20 - 12a - 1 \\ 3a - 10a + 12a = 20 + 2 - 1 - 11 \\ (3 - 10 + 12)a = 22 - 12 \\ (15 - 10)a = 10 \\ 5a = 10 \\ a = \frac{10}{5} \\ a = 2 \)